OpenStudy (anonymous):

Simplify (5^2-3x+)-(x^2+2x-1)

OpenStudy (anonymous):

you should distribute the negative one first

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

help

OpenStudy (anonymous):

so you see a negative in front of a parenthesis right? that means you can multiply everything inside the parenthesis by -1 can you do that part for me?

OpenStudy (anonymous):

know

OpenStudy (anonymous):

i show you \[-(x^2 + 2x - 1) = -x^2 - 2x + 1\]

OpenStudy (anonymous):

so now you need to simplify this \[5^2 - 3x - x^2 - 2x + 1\] just wondering is it 5^2 or 5x^2?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

im already trying to help you im not gonna just give you the answer you know

OpenStudy (anonymous):

is it 5x^2

OpenStudy (anonymous):

i was asking you if you made a typo or not in the original problem cause you wrote 5^2

OpenStudy (anonymous):

if you didnt make a typo then i'll keep going

OpenStudy (anonymous):

no I did not

OpenStudy (anonymous):

okay tell me whats 5^2?

OpenStudy (anonymous):

and tell me whats -3x - 2x?

OpenStudy (anonymous):

i dot know

OpenStudy (anonymous):

5^2 = 25 and -3x - 2x = -5x so your answer is -x^2 - 5x + 26

OpenStudy (anonymous):

yeah i got the same thing as @jayz657 @bigpapa021201

OpenStudy (anonymous):